sam is 10 years older than jack.twice sam`s age subtracted from 6 times jack`s age is 7 more than jack`s age. find jack and sam`s age
Answer by ankor@dixie-net.com(12684)

s = sam's age now; j = jacks age now
:
Write an equation for each statement:
:
"sam is 10 years older than jack."
s = j + 10
:
"twice sam`s age subtracted from 6 times jack`s age is 7 more than jack`s age."
6j - 2s = j + 7
6j - j - 2s = 7; subtract j from both sides
5j - 2s = 7
Substitute (j+10) for s (the 1st equation)
5j - 2(j+10) = 7
5j - 2j - 20 = 7
3j = 7 + 20
3j = 27
j = 27/3
j = 9 yrs old
:
obviously s = 19 yrs old
:
:
Check solutions using the statement:
"twice sam`s age subtracted from 6 times jack`s age is 7 more than jack`s age."
6(9) - 2(19) = 9 + 7
54 - 38 = 16